Using elimination of arbitrary constant
y^2 (a-x) = x^2 (a+x)
1
Expert's answer
2020-10-13T19:17:29-0400

Solution:


"y^2 (a-x) = x^2 (a+x)"

Now, take derivative w.r.t x on both side we get,

"2yy'(a-x)-y^2=2ax+3x^2\\
\implies a = \frac{2yxy'-3x^2-y^2}{2yy'+2x}"

Now, substitute a in original equation we get,


"y^2(-5x^2-y^2)=x^2(4yxy'-x^2-y^2)\\"